OpenStudy (anonymous):

What is the y-intercept of the equation? y = 2.1(0.15)^x a.) 0.15 b.) 1 c.) 2.1

OpenStudy (tony123):

plug in 0 for x to get y= 2.1 (0.15)^0 anything to the power of 0 =1 so we get 2.1 *1= 2.1 C